OpenStudy (anonymous):

A team of 3 students is to be formed from a class of 25 students. How many ways are there to form this team if there is 1 team leader, 1 scribe and 1 timekeeper in the team and each person can only perform one role at a time?

OpenStudy (anonymous):

\[25\times 24\times 23\]

OpenStudy (anonymous):

May i know how did you get it?

OpenStudy (anonymous):

called "counting principle" if there are n ways to do one thing and n to do the other then there are \[m\times n\] ways to do them both

OpenStudy (anonymous):

so reason as follows. you have 25 choices for "team leader". once team leader is chosen you then have 24 choices for "scribe" (who writes this stuff?) so far 25*24 and once they are chosen you have 23 ways to pick timekeeper. so answer is as i wrote above

OpenStudy (anonymous):

oh thanks :)